simplify and solve the system using elimination
2(x-4) +3(y+2) =-8
4(x+1) +5(y-1)= -9

Answer by Edwin McCravy(20055)

Simply the first equation:

2(x-4)+3(y+2) = -8
    2x-8+3y+6 = -8
      2x+3y-2 = -8
        2x+3y = -6

Simplify the second equation:

4(x+1)+5(y-1) = -9
    4x+4+5y-5 = -9
      4x+5y-1 = -9
        4x+5y = -8

Solve the simplified system:

        2x+3y = -6
        4x+5y = -8

To eliminate x, multiply the first equation by -2
to cause the x terms to cancel when adding them term
by term:

       -4x-6y = 12
        4x+5y = -8
       -----------
           -y = 4
            y = -4 

Substitute -4 for y in

        2x+3y = -6
     2x+3(-4) = -6
        2x-12 = -6
           2x = 6
            x = 3

Solution: (x,y) = (3,-4)

Edwin